FEL Upgrade Project Weekly Brief
October 28 - November 1, 2002

Zone 4 (the second cryomodule) is ready for checkout with RF and srf commissioning to full gradient. The extraction region dipole magnets (to the energy recovery dump) were installed and aligned completing all iron magnet installation required for linac operation to the straight ahead dump. The vacuum chamber has passed checkout and is ready to go into them on Monday. The gun electrode assembly was completed and undergoing final check before installation in the tunnel early next week. The UV optical cans arrived on site.

A cost and schedule review was held with G. Graff and John Albertine to discuss closeout of the FY02 construction budget, FY03 budget issues, and scaleup plans to higher power.

JLab hosted IPF2002 on Oct. 27-29 with 70 attendees from academia, industries, and the government. There were interesting presentations on a variety of physics and societal related topics as well as tours of CEBAF and the FEL.

The pi bend vacuum chamber was received from the vendor and is being prepared for leak and dimensional checks.

We completed the September Monthly report.

G. Neil was interviewed by the Science Office of the EPSRC of the UK as part of the Gateway review process of the 4GLS Project at Daresbury Laboratory. They were presenting the "Business Case" to go get permission to proceed with full design and component development. (similar to CD-1 in DOE-speak).

Several papers were approved for presentation at the DEPS conference later this month.

WBS 4 (Injector):
Cesium channels were installed in the ball and final assembly of the stack is being completed with the door assembly requiring a little more adjustment. The external corona rings and rods were checked/assembled for initial installation onto the stack. The spacer flange for the stack was readied for installation. An NF3 reservoir was mounted onto the gun stand. The solenoids fields were measured for proper polarity.

WBS 6 (RF):
Zone 4 - All interlocks have been checked. The helium pressure transducer will be replaced next week. SRF Commissioning will start today.

Zone 2 - This zone is ready for SRF testing also.

Buncher - In order to have a 250°C bake of the Gun and adjacent beamline, the fittings will be removed from the Buncher. Since there is a mechanical interference between 1 right-angle water fitting and the Buncher mounting bracket, the filling will be cut out and replaced with an in-line fitting. The o-rings of all fittings will be replaced with silicon rubber o-rings.

Quarter HVPS - No progress this week, concentrated on zone 4.

Gun HVPS - Additional parts were designed, fabricated, and installed to ensure a good contact through the Running/Conditioning Resistor mechanism. The control box to operate the resistor mechanism will be finished early next week. Resistors were added to the Glassman HVPS to enable operation to 600 kV. It will be tested and calibrated next week.

WBS 8 (Instrumentation):
Drive Laser Controls - The new 8-Ch. stepper motor controller has checked-out well and the variable attenuator is working from the expert screens in EPICS. Al Grippo will set-up the top-level commands next week. The MPS shutter has a similar status. The hardware and low-level software checks-out, but the high level EPICS commands must be modified to reflect the changes. Work has begun to make the external RF interlock board that will switch the drive laser to a local MO to keep the drive laser mode-locked in the event of a drop-out of external RF. This board will also carry a single axis picomotor and a small micro-controller to provide for closed loop control of the drive laser's optical cavity based on the phase error from the low-level RF control module. This upgrade will save us from having to make routine cavity length adjustments on the drive laser. A description can be found in the white paper listed at the following link:


Drive Laser Pulse Controller (DLPC) - The carrier board for the Conoptics Model 305 synchronous countdown chassis has been assembled and is being tested. As the machine protection functions of the DLPC are to be disarmed in the case where the FEL injector's high-voltage power supply interface is determined to be OFF, the necessary cabling is in progress as well as the logic to fiber module which will be implemented into iocfel10.

The Machine Protection System (MPS) is progressing well. The initial programming is complete for the Beam Mode and Machine Mode input boards. Front panels are being manufactured. The fiber input (Fast Shut Down FSD) & output (to DLPC) board is in the final stage of layout. A few questions come up regarding details of the various modes, these will be resolved early next week.

The skew quad cabling from the 0F to the 3F regions have been routed and terminated. Control screens for corrector magnets in the 1F and 2F regions and the Extraction dipole have been completed and waiting for checkout. Field measurements were made on solenoid magnets MFF0F01 and MFF0F02 to test for polarity. We spent some time to work on the magnet measurement test stand. There are a number of cabling issues that have not been properly addressed for more than 10 years. A number of the resolver and signal cables were re-terminated which solved the bulk of the intermittent problems that has been slowing down the throughput.

Final alignment/ testing of the Beamviewers and cameras in the 0F to the 3F regions has been completed. The Dump Skid chassis has been installed. Bench checkout of the HV interlock chassis is continuing. Parts have been ordered to complete the Beamviewer Camera Mounting Assemblies. New camera boxes were machined, however corrections need to be made on the dimensions of the holes that were drilled. Additional lamp/interface boxes are being assembled as are additional cameras. Cabling for the Optical Cavity (Output Coupler and High Reflector) have been installed for the 5F region. The interface chassis that converts the 12 pin motor drive connectors to the 25 pin connectors for the optical cavity vacuum feedthroughs is complete.

The EECad group has recently provided the majority of our back-logged circuit board projects. Thanks Very Much!!! As a result the check-prints and notes to revise and archive are beginning to pile up .

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
The Purcell Gap sheet pieces are starting to be attached to the tops of the six magnets after the field clamps were pinned and the tooling ball holes drilled.
• The comparison test stand problem was solved by Kevin Jordan and his crew by changing out some cables that had bad terminations.
• An additional GW was placed on the stepper stand and started field clamp configuration for final measurements. The measurement technicians have to learn new techniques to adjust these large field clamps. After some preliminary adjustment tries, they ordered some simple custom tools made through the machine shop to make the adjustment process efficient.
Injector Dipoles (DU/DV)
• We again machined the pole face ends on the GUs because the first machining pass brought us just outside of the range that allows field clamp adjustment to bring the magnet to specified field integral. After a final measurement, the magnet was released, installed and aligned along with the companion GVs in the extraction line. This completes the magnet installation in the Linac for first commissioning.
• The last GU was measured and found to have an anomaly in the core field through the central cross section It has a warp to a portion of the Purcell gaps of the upper pole tip. We will probably have to strip these sheets off and re glue them. These were measured and passed by the vendor (Wang NMR) and so failed since we received them, perhaps during the edge machining.
Arc 180 Degree Dipoles (GY)
• Wang NMR has stepped up their work on coils to produce one GY by the end of November after getting the GQs and GXs ready for shipment.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• The first two units of each style of magnet are ready for shipment, awaiting a shipper’s truck.
• Work continued in potting the remaining 4 GX coils for the next shipment.
Trim Quad (QT)
• No additional progress to report.
Sextupole (SF)
• Master Machine continues making the shim parts. They will receive the original SCs next Tuesday for disassembly.
Octupole (OT)
• We continue working on a manufacturing package.
Beam Line and Vacuum
• Master Machine continues in production status where one arc chamber a week is coming off the production line.
• Injection line had all magnets after the quarter Cryomodule installed
• Chicane chambers continue fit up, welding and machining

WBS 11 (Optics):
Visitors to User Lab 2 would immediately think, "new car smell", from all the extra plastic sheeting hung to increase the clean area. Since last week's report, one of the "top hats" of the optical cavity vacuum vessels has the parts welded to it, and it has been cleaned for UHV use. We will leak check it here at the FEL this afternoon. The second top hat will be worked on later today and early next week. All optical cavity internals were cleaned for UHV, and are mostly assembled in the clean environment. We discovered during the assembly that some of the Dicronited parts (a dry lubricant process) were stripped clean and a few other parts require silver plating before we can put the optical vessels under vacuum. After some discussions with the Survey and Alignment group, we decided that the parts can be replaced without jeopardizing the alignment and fiducialization process. This is rather important, because they are available to do our work for only about another week before heading over to CEBAF to catch up on some work there. Survey and Alignment began their work by gluing alignment targets to the outside of the vacuum vessels.

The outcouplers for first light, and high power operation at ~ 10 microns arrived from the vendor ~ 1 week ahead of schedule. They are being inspected. Assuming they pass inspection, we will have the optics required for first light. The ultraviewers that were assembled last week were aligned and are ready for integration onto the support girder. Installation of the first alignment HeNe for the optical hardware (and later, the wiggler and beamline viewers) began yesterday.

Drawings for the optical transport turning mirror vacuum vessels will be signed off today. The collection mirror for the 0F region THz diagnostic was installed and is being aligned.

We received the vacuum vessels for the UV Upgrade FEL.

Other activities:
We participated in the Industrial Physics Forum tour.

Commissioning/Operations Activities:
We checked the calibration of the drive laser IR monitor.